OpenStudy (anonymous):

Every evening Jenna empies her pockets and puts her change in a jar. At the end of the week she counts her money. One week she had 40 coins, all of them dimes and quarters. When she added them up, she had a total of $7.75. d= number of dimes q= number of quarters How many quarters did Jenna have?

OpenStudy (anonymous):

d+q=40 but I don't know what to do next

OpenStudy (anonymous):

\[( d + q = 40, 10 d + 25 q = 775 )\]

OpenStudy (anonymous):

The second equation concerns pennies. { d -> 15, q -> 25 }

OpenStudy (anonymous):

ok so then you can substitute the values and find the answer

OpenStudy (anonymous):

so 15+25=40

OpenStudy (anonymous):

I got 15 dimes and 25 quarters

OpenStudy (anonymous):

Replace d below with 40-q 10 d + 25 q = 775 /. d -> 40 - q solve the following for q: 10 (40 - q) + 25 q = 775

OpenStudy (anonymous):

15*10+25*25=775

OpenStudy (anonymous):

thank u :D